A hundred grams of Sulfuric acid yielded ten grams of water.

H2SO4(l) = H2O(l) + SO3(l)


a. Balance the chemical equation;

b. Identify the limiting reactant and the excess reactant if applicable;

c. Compute for the theoretical yield; d. Determine the percent yield of the reaction;

e. Calculate the percent error and

f. Compute for the excess amount of the excess reactant if applicable.

g. Show the complete solutions for your answers in a separate paper.


1
Expert's answer
2021-04-14T06:09:45-0400

Balanced

H2SO4(l) = H2O(l) + SO3(l)


Moles of H2SO4 = "\frac{100}{98}=1.02 moles"


Moles of H2O "=1.02 moles"


Not applicable for limiting and excess reactant


Theoretical yield = 1.02 × 18=18.36 g


Percent yield = "\frac{18}{18.36}×100=98" %